### v2

@@ IV.3 Results when $\gamma=1$
The situation considered when $\gamma=1$ was studied in detail in Ref. 1 and the property of locality proved in Ref. 15. However, for the sake of completeness, we show some results for $\gamma=1$ which are the counterparts of those illustrated in Fig. 4 and Figs. 6 to 8.
@@ IV.3.1 $A=1$, $v_{0}=50$
As may be clearly see in Fig. 8, when $\gamma=1$, we do find an initial quasilinear regime when $A=1$ and $v_{0}=50$ and the time evolution of $\langle\Delta v^{2}(t)\rangle$ looks essentially independent of $N$, up to some fluctuations due to the finite number of phase realizations used in our averages.
@@
As already mentioned in Paragraph IV.1, this was interpreted in Ref. 1 as the direct consequence of locality leading to the crossover between the initial regime of initial quasilinear diffusion and the later regime of chaotic diffusion.
@@ IV.4.1 $A=1$, $v_{0}=50$
Figure 11: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=0.5$, $A=1$ and $v_{0}=50$. The red dashed solid line is for $N=109$, the black solid line for $N=685$.
@@
Figure 12: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=0.5$, $A=11.5$ and $v_{0}=100$. The red dashed solid line is for $N=218$, the black solid line for $N=1374$.
@@ V Chaotic transport modeled as a diffusion process limited by phase space boundaries
In this section, we introduce a heuristic simple formula to derive the time evolution of the velocity distribution function in the situation when there is bounded chaotic transport, as defined in Section II. We restrict to regimes corresponding Figs. 10 and 12, i.e., a quasilinear diffusion as long as the orbits remain far enough from the phase space boundaries. However, we aim to derive the time evolution of the distribution function for arbitrarily long times. This is not a simple task since the exact equation we have to solve is actually unknown. In particular, although Dirichlet boundary conditions seem most appropriate for this problem, solving the diffusion equation with such boundary conditions would not yield the correct answer because $\int f(v)dv$ would not be conserved. Hence, we only propose here a heuristic formula for the time evolution of the distribution function, that we compare against results from test particles simulations. Clearly, this is a preliminary study that would require a more in-depth investigation, to better understand the range of validity of our formula and to address the more general situation when the diffusion coefficient is not uniform. However, we found useful to include our formula in this article due to its very good agreement with the numerical results. Hopefully, it will help improve the modeling of bounded chaotic transport.
@@
As for $v_{\min}$ and $v_{\max}$ they are estimated the following way. Let $n_{\min}$ and $n_{\max}$ be such that $\min(\omega_{n}/k_{n})=\omega_{n_{\min}}/k_{n_{\min}}$ and $\max(\omega_{n}/k_{n})=\omega_{n_{\max}}/k_{n_{\max}}$. Then, in Eq. (60), we use the following expressions for $v_{\min}$ and $v_{\max}$,
@@
Fig. 17 compares results obtained from Eq. (60) against those derived from test particles simulations. We use for $f(v,0)$ a Maxwellian, $f(v,0)=e^{-v^{2}/2}/\sqrt{2\pi}$ when $v<2$ and $v>4$, and a symmetrized Maxwellian, $f(v,0)=\left[e^{-v^{2}/2}+e^{-(2v-3)^{2}/2}\right]/2\sqrt{2\pi}$ when $|v-3|\leq 1$. $f_{0}(v)\equiv f(v,0)$ is illustrated in Fig. 16. This is typically the distribution function resulting from the adiabatic trapping of electrons in a plasma wave [23, 22]. Such a wave may be subjected to instabilities leading to the growth of sidebands [24]. Our work was motivated by the evolution of the electron distribution function under the action of these sidebands, whence the choice $f(v,0)$. As for the test particle simulations, we numerically solve the Hamilton equations of,
@@
Fig. 17 shows that Eq. (60) provides results in very good agreement with those from test particle simulations except close to $v=2$ when $D_{\rm{QL}}t\lesssim 0.5$. These slight discrepancies are expected because $f_{0}(v)$ is discontinuous at $v=2$ and a diffusion equation is not valid about a discontinuity for short times. This is easily proven the following way. Let us a consider a distribution function, $f_{0}(v)$, discontinuous at $v=v_{d}$, and let $\Delta f_{0}=f_{0}(v_{d}^{-})-f_{0}(v_{d}^{+})$ be the jump in $f_{0}$ at the discontinuity. Cleary, for any particle and at any time, the change in velocity induced by the wave spectrum remains finite. Consequently, the particles which have crossed the discontinuity by time $t$ have initial velocities, $v_{0}$, such that $|v_{0}-v_{d}|\leq\delta v_{\max}(t)$, where $\delta v_{\max}(t)$ is a finite velocity increment that grows with time and such that $\delta v_{\max}(t=0)=0$. Then, at the discontinuity and at time $t$, the change in $f_{0}$ induced by the wave spectrum is
@@
where $\eta(v_{0})$ is the fraction of particles with initial velocity, $v_{0}$, which have crossed the discontinuity. Clearly, $\delta f(0)=0$ and, although $\delta f(t)$ is not necessarily a monotonic function of time, it would clearly tend to increase with $t$. Consequently, it requires a finite time for $\delta f$ to be larger than the initial jump in $f_{0}$, $\Delta f_{0}$. Before this time, the distribution function remains discontinuous. This is contrast with the solution of a diffusion equation which becomes differentiable whenever $t>0$.
@@ VI Conclusion
This paper has provided a simple criterion for assessing its validity, which is a central issue in many areas of physics and, most notably, in plasma physics. For a smooth discrete wave spectrum, the perturbative regime of quasilinear diffusion, akin to that introduced in the original 1962 papers, is valid only when the series on the right-hand side of Eq. (41) converges. This condition is equivalent to the requirement of locality in phase velocity for the wave-particle interaction. When this criterion is satisfied then, as illustrated in Figs. 14 and 15 of Paragraph IV.5, one can define a broad class of Hamiltonians whose transport properties are exactly the same as those of Hamiltonian $H_{p}$ [Eq.(1)]. When the condition for locality is not fulfilled, the early time evolution of the velocity spreading is not diffusive, as clearly shown in Fig. 4 of Paragraph IV.1.1. Moreover, it depends on the number of modes in the discrete spectrum, unlike what quasilinear theory would predict. In spite of these very clear defects we have to acknowledge that, in all our simulations, quasilinear theory does provide the right order of magnitude for $\langle\Delta v^{2}(t)\rangle/t$ at short times, with discrepancies never exceeding a factor of about two.

### v3

@@ IV.3 Results when $\gamma=1$
The situation considered when $\gamma=1$ was studied in detail in Ref. 1 and the property of locality proved in Ref. 16. However, for the sake of completeness, we show some results for $\gamma=1$ which are the counterparts of those illustrated in Fig. 4 and Figs. 6 to 8.
@@ IV.3.1 $A=1$, $v_{0}=50$
As may be clearly seen in Fig. 9, when $\gamma=1$, we do find an initial quasilinear regime when $A=1$ and $v_{0}=50$ and the time evolution of $\langle\Delta v^{2}(t)\rangle$ looks essentially independent of $N$, up to some fluctuations due to the finite number of phase realizations used in our averages.
@@
As already mentioned in Paragraph IV.1, this was interpreted in Ref. 1 as the direct consequence of locality leading to the crossover between the initial regime of quasilinear diffusion and the later regime of chaotic diffusion.
@@ IV.4.1 $A=1$, $v_{0}=50$
Figure 11: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=0.5$, $A=1$ and $v_{0}=50$. The red dashed solid line is for $N=109$, the black solid line for $N=687$. These values for NN correspond to a maximum phase velocity being, respectively, 2 and 20 times $v_{0}$.
@@
Figure 12: Time dependence of $\langle\Delta v^{2}(t)\rangle/2D_{\rm{QL}}t=\langle\Delta v^{2}(t)\rangle/\pi A^{2}t$ when $\gamma=0.5$, $A=11.5$ and $v_{0}=100$. The red dashed solid line is for $N=218$, the black solid line for $N=1374$. These values for NN correspond to a maximum phase velocity being, respectively, 2 and 200 times $v_{0}$.
@@ V Chaotic transport modeled as a diffusion process limited by phase space boundaries
In this section, we introduce a heuristic simple formula to derive the time evolution of the velocity distribution function in the situation when there is bounded chaotic transport, as defined in Section II. We restrict to regimes corresponding to Figs. 10 and 12, i.e., a quasilinear diffusion as long as the orbits remain far enough from the phase space boundaries. However, we aim to derive the time evolution of the distribution function for arbitrarily long times. This is not a simple task since the exact equation we have to solve is actually unknown [23]. In particular, although Dirichlet boundary conditions seem most appropriate for this problem, solving the diffusion equation with such boundary conditions would not yield the correct answer because $\int f(v)dv$ would not be conserved. Hence, we only propose here a heuristic formula for the time evolution of the distribution function, that we compare against results from test particles simulations. Clearly, this is a preliminary study that would require a more in-depth investigation, to better understand the range of validity of our formula and to address the more general situation when the diffusion coefficient is not uniform. However, we found useful to include our formula in this article due to its very good agreement with the numerical results. Hopefully, it will help improve the modeling of bounded chaotic transport.
@@
Note that, when $\sqrt{2D_{\rm{QL}}t}\ll(v_{\max}-v_{\min})$ and when $v_{\max}-v\gg\sqrt{2D_{\rm{QL}}t}$ and $v-v_{\min}\gg\sqrt{2D_{\rm{QL}}t}$, the integral in the denominator of the right-hand side of Eq. (60) is close to $\sqrt{4\pi D_{\rm{QL}}t}$ so that Eq. (59) is recovered, as should be. When $\sqrt{2D_{\rm{QL}}t}\gg(v_{\max}-v_{\min})$, Eq. (60) predicts a nearly constant $f(v)$ within the interval $v_{\min}\leq v\leq v_{\max}$, as expected and as may be appreciated in Fig. 17. As for $v_{\min}$ and $v_{\max}$ they are estimated the following way. Let $n_{\min}$ and $n_{\max}$ be such that $\min(\omega_{n}/k_{n})=\omega_{n_{\min}}/k_{n_{\min}}$ and $\max(\omega_{n}/k_{n})=\omega_{n_{\max}}/k_{n_{\max}}$. Then, in Eq. (60), we use the following expressions for $v_{\min}$ and $v_{\max}$,
@@
Fig. 17 compares results obtained from Eq. (60) against those derived from test particles simulations. We use for $f(v,0)$ a Maxwellian, $f(v,0)=e^{-v^{2}/2}/\sqrt{2\pi}$ when $v<2$ and $v>4$, and a symmetrized Maxwellian, $f(v,0)=\left[e^{-v^{2}/2}+e^{-(6-v)^{2}/2}\right]/2\sqrt{2\pi}$ when $|v-3|\leq 1$. $f_{0}(v)\equiv f(v,0)$ is illustrated in Fig. 16. This is typically the distribution function resulting from the adiabatic trapping of electrons in a plasma wave [25, 24]. Such a wave may be subjected to instabilities leading to the growth of sidebands [26]. Our work was motivated by the evolution of the electron distribution function under the action of these sidebands, whence the choice $f(v,0)$. As for the test particle simulations, we numerically solve the Hamilton equations of,
@@
Fig. 17 shows that Eq. (60) provides results in very good agreement with those from test particle simulations except close to $v=2$ when $D_{\rm{QL}}t\lesssim 0.5$. These slight discrepancies are expected because $f_{0}(v)$ is discontinuous at $v=2$ and a diffusion equation is not valid about a discontinuity for short times. This is easily proved the following way. Let us a consider a distribution function, $f_{0}(v)$, discontinuous at $v=v_{d}$, and let $\Delta f_{0}=f_{0}(v_{d}^{-})-f_{0}(v_{d}^{+})$ be the jump in $f_{0}$ at the discontinuity. Cleary, for any particle and at any time, the change in velocity induced by the wave spectrum remains finite. Consequently, the particles which have crossed the discontinuity by time $t$ have initial velocities, $v_{0}$, such that $|v_{0}-v_{d}|\leq\delta v_{\max}(t)$, where $\delta v_{\max}(t)$ is a finite velocity increment that grows with time and such that $\delta v_{\max}(t=0)=0$. Then, at the discontinuity and at time $t$, the change in $f_{0}$ induced by the wave spectrum is
@@
where $\eta(v_{0})$ is the fraction of particles with initial velocity, $v_{0}$, which have crossed the discontinuity. Clearly, $\delta f(0)=0$ and, although $\delta f(t)$ is not necessarily a monotonic function of time, it would clearly tend to increase with $t$. Consequently, it requires a finite time for $\delta f$ to be larger than the initial jump in $f_{0}$, $\Delta f_{0}$. Before this time, the distribution function remains discontinuous. This is in contrast with the solution of a diffusion equation which becomes differentiable whenever $t>0$.
@@ VI Conclusion
This paper has provided a simple criterion for assessing its validity, which is a central issue in many areas of physics and, most notably, in plasma physics. For a smooth discrete wave spectrum, the perturbative regime of quasilinear diffusion, akin to that introduced in the original 1962 papers, is valid only when the series on the right-hand side of Eq. (41) converges. This condition is equivalent to the requirement of locality in phase velocity for the wave-particle interaction. When this criterion is satisfied then, as illustrated in Figs. 14 and 15 of Paragraph IV.5, one can define a broad class of Hamiltonians whose transport properties are very similar to those of Hamiltonian $H_{p}$ [Eq.(1)]. When the condition for locality is not fulfilled, the early time evolution of the velocity spreading is not diffusive, as clearly shown in Fig. 4 of Paragraph IV.1.1. Moreover, it depends on the number of modes in the discrete spectrum, unlike what quasilinear theory would predict. In spite of these very clear defects we have to acknowledge that, in all our simulations, quasilinear theory does provide the right order of magnitude for $\langle\Delta v^{2}(t)\rangle/t$ at short times, with discrepancies never exceeding a factor of about two.
@@
Acknowledgements.
The author acknowledges the very careful reading and comments by D. Escande and the useful discussions with Y. Elskens.